OpenStudy (calculusxy):

The NY Knicks spent $2610 on equipment at the end of last season. Basketball costs $45 and new nets for the basket cost $15. If the Knicks have 66 new balls and nets combined this season, how many basketballs did they buy? How many nets did they buy?

OpenStudy (calculusxy):

@Mertsj @SolomonZelman

OpenStudy (anonymous):

B= basketballs N=Nets $45b+ $15n= 2610 b==66 n=?

OpenStudy (anonymous):

How many nets did they buy?

OpenStudy (calculusxy):

It's saying that 66 equals to the total number of balls and nets. x = # of basketballs y = # of nets x + y = 66

OpenStudy (anonymous):

b + n = 66 $45b+ $15n= 2610

OpenStudy (anonymous):

we can solve now

OpenStudy (calculusxy):

So let's say we want to figure out the number of basketballs. $45b + $15n = $2610 $45b + $15n - $15n = $2610 - $15n $45b = $2610 - $15n

OpenStudy (calculusxy):

@welshfella @DanJS

OpenStudy (anonymous):

\[\frac{ 1 }{45 }($45b = $2610 - $15n)\]

OpenStudy (calculusxy):

I don't understand that.

OpenStudy (anonymous):

b=58-1/3n

OpenStudy (anonymous):

b+n=66 b= 58-1/3n 58-1/3n+n=66 (2/3)n= 8 n=12

OpenStudy (anonymous):

b+12=66 b=54

OpenStudy (anonymous):

Do you get it now?

OpenStudy (calculusxy):

Yes thank you